Given that X is a continuous random variable that has a uniform probability distribution, and 0...

  1. Given that X is a continuous random variable that has a uniform probability distribution, and 0 < X < 8:

a. Calculate P(X < 4) (to 3 significant digits).
P(X < 4)=

b. Determine the mean (µ) and standard deviation (σ) of the distribution (to 3 significant digits).
µ =


σ =
Answer #1

Solution :

Given that,

a = 0

b = 8

(a)

P(x < 4) = (4 - 0) / (8 - 0) = 0.5

(b)

mean = µ = (a + b) / 2 = (0 + 8) / 2 = 4

σ = [(b - a)2 / 12] = [(8 - 0)2 / 12] = 2.309